# Inductive Gradient Adjustment for Spectral bias in implicit neural representations

### v1

@@ 5 Empirical Analysis on Simple Function Approximation
In this experiment, we firstly give an empirical validation to our Theorem 4.1 that impacts of NTK-based and eNTK-based adjustments on spectral bias are similar. Then, we demonstrate that our inductive gradient adjustment (IGA) method can purposefully improve the spectral bias of MLPs by managing the number of balanced eigenvalues of the empirical inductive NTK matrix $\tilde{{\bm{K}}}_{e}$.
To better analyze, we follow previous works (Xu, 2018; Shi et al., 2024a) and compute the relative error $\Delta_{k}$ between target signal $g$ and outputs $f_{\Theta}$ at frequency $k$ to show spectral bias:
$\Delta_{k}=\frac{|\mathcal{F}_{D}[g](k)-\mathcal{F}_{D}[f_{\Theta}](k)|}{|\mathcal{F}_{D}[g](k)|},$ (7)
@@
Experiment 1. In this experiment, we aim to corroborate our theoretical results from the previous section, specifically: impacts of eNTK-based gradient adjustment on spectral bias are similar to that of NTK-based gradient adjustment, and the differences decrease as the network width increases; inductive generalization via $\tilde{{\bm{K}}}_{e}$ for gradient adjustment is effective, and increasing width helps reduce the estimate error. Therefore, the accurate computation of NTK matrix ${\bm{K}}$ is the key for validation. Given that the accurate computation of ${\bm{K}}$ is non-trivial, we adopt a two-layer MLP with a fixed last layer in Arora et al. (2019); Ronen et al. (2019), whose ${\bm{K}}_{ij}$ can be computed by the formula $\frac{1}{4\pi}(\mathbf{x}_{i}^{\top}\mathbf{x}_{j}+1)(\pi-arccos(\mathbf{x}_{i}^{\top}\mathbf{x}_{j}))$. This architecture facilitates theoretical analysis but has limited representational capacity for representing complex functions. To better demonstrate the spectral bias of MLPs, we constructed the following simple function $f:\mathbb{S}^{1}\rightarrow\mathbb{R}^{1}$ that has been widely adopted by previous works (Ronen et al., 2019; Geifman et al., 2023):
$f(\theta)=f(cos(\theta),sin(\theta))=sin(0.4\pi\theta)+sin(0.8\pi\theta)+sin(1.6\pi\theta)+sin(3.2\pi\theta).$ (8)
@@
Then the two-layer MLP in Arora et al. (2019) is employed to parameterize this function with $N=1024$ input samples by sampling $\theta$ in $[0,2\pi]$. We vary the width of the MLP from $1024$ to $8192$. We adopt the Identity matrix ${\bm{I}}$ (i.e., vanilla gradient) and a series of ${\bm{S}},\tilde{{\bm{S}}},\tilde{{\bm{S}}}_{e}$ that $start=1$ and $end$ ranges from $11$ to $15$. The groups of our IGA method are partitioned by the order of $\theta$ and the size is $8$, indicating that $\tilde{{\bm{S}}}_{e}$ is only $1/64$ the size of ${\bm{S}},\tilde{{\bm{S}}}$. All MLPs are trained separately with the same fixed learning rates by SGD for $20000$ iterations. All models are repeated 10 times with different random seeds. We visualize results of baseline model and adjustments with $end=15$ in Fig. 1. More results can be found in our Appendix C.
Figure 1: Evolution of approximation error with training iterations on time domain and Fourier domain. Line plots visualize the MSE loss curves of MLPs with $1024$ and $4096$ neurons optimized by four methods. The shaded area of lines indicates training fluctuations, represented by twice the standard deviation. Heatmaps show the relative error $\Delta_{k}$ on four frequency bands. More results can be found in our Appendix C.
@@
Analysis of Experiment 1. In Fig. 1, line plots on a gray background show the convergence trends of MLPs by NTK-based (${\bm{S}}$), eNTK-based ($\tilde{{\bm{S}}}$) and NTK-based ($\tilde{{\bm{S}}}_{e}$) gradient adjustments. Vanilla gradient (${\bm{I}}$) is also compared. The heat maps with a consistent color scale visualize their impacts on the spectrum, where darker colors indicate higher errors. Note that initial color bars of vanilla gradients (${\bm{I}}$) darken rapidly from $0.4$Hz as frequency increases. This means that the MLP optimized by vanilla gradient suffers severe spectral bias. With gradient adjustments, there are more lighter shades over frequencies $0.4$Hz and $0.8$Hz, indicating that these adjustments effectively improve the spectral bias. Therefore, MLPs with gradient adjustments have faster convergence rate in MSE loss as shown in line plots. Please note that although adjustments by ${\bm{S}}$, $\tilde{{\bm{S}}}$ and $\tilde{{\bm{S}}}_{e}$ exhibit slight difference in trends and impacts on spectrum when the network width is 1024, the difference is barely noticeable when the width is 4096. This is consistent with our Theorem 4.1 and the analysis of our Theorem 4.2. Moreover, despite the presence of difference when the network width is 1024, $\tilde{{\bm{S}}}$ and $\tilde{{\bm{S}}}_{e}$ effectively improve spectral bias and the approximation accuracy, indicating that $\tilde{{\bm{K}}}$ successfully links spectral bias with training dynamics and our IGA method is effective.
@@
Experiment 2. In this experiment, we aim to show that our IGA method can tailor impacts on spectral bias of general architectures by managing the number of balanced eigenvalues and analyze tailored impacts under varying group sizes. Therefore, we apply our method on two practical INR models, i.e., MLPs with ReLU (ReLU) and MLPs with Sin (SIREN) (Sitzmann et al., 2020) like Shi et al. (2024a). To analyze impacts under varying sizes, we draw on settings in Rahaman et al. (2019) and construct a 1D function $f:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$ with abundant spectrum as follows:
$f(x)=sin(20\pi x)+sin(40\pi x)+sin(60\pi x)+sin(80\pi x)+sin(100\pi x)+sin(120\pi x).$ (9)
@@
SIREN is trained to regress the $f(x)$ with $2048$ discrete values uniformly sampled in the interval $[-1,1]$. For ReLU, we halve the frequency of each component due to its limited representation capacity (Yüce et al., 2022). For our IGA method, we partition data points using two interval lengths of $4$ and $8$ to compute $\tilde{{\bm{K}}}_{e}$ and varies $end$ from $2$ to $8$ to construct a series of $\tilde{{\bm{S}}}_{e}$. ENTK-based gradient adjustment is considered. All gradient adjustments are conducted across multiple MLPs with a four-hidden-layer, 256-width architecture, optimized by Adam for 10,000 iterations with a fixed learning rate of $55e-$ 5. All models are repeated 10 times with different random seeds.
Figure 2: Progressively amplified impacts on spectral bias of ReLU and SIREN by increasing the number of balanced eigenvalues of $\tilde{{\bm{S}}}_{e}$ when the group size is $8$. ReLU denotes that the MLP with ReLU optimized using vanilla gradients; ReLU-$\tilde{{\bm{S}}}_{e}$-2 denotes that the MLP with ReLU optimized using gradients adjusted by $\tilde{{\bm{S}}}_{e}$ with $end=2$. More results can be found in our Appendix C.
Figure 3: Comparison of group sizes with varying balanced eigenvalues on Relative Error at 20Hz, 40 Hz. IGA-1 denotes that the group size is 1, i.e., the eNTK-based gradient adjustment; IGA-4 denotes that the group size is 4, i.e., inductive gradient adjustment; Baseline denotes that the MLP is optimized by vanilla gradients. The shaded area of lines indicates training fluctuations, represented by twice the standard deviation. Results on other frequencies can be found in Appendix C.
@@
Analysis of Experiment 2. In Fig. 2, with more balanced eigenvalues, the proportion of lighter regions in heatmaps of both ReLU and SIREN grows, while the curve of relative error decreases more rapidly in Fig. 3. These observations clearly illustrate that impacts on spectral bias by our IGA method can be amplified by increasing the number of balanced eigenvalues. Further, as shown in Fig. 3, despite the relative error curves of ReLU exhibit subtle differences as the group size increases, our IGA method effectively amplifies impacts on spectral bias by increasing the number of balanced eigenvalues.
@@ 6 Experiment on Vision Applications
In this section, we apply our IGA method to various practical applications of INRs in computer vision, demonstrating the superiority of IGA.
@@
It is well known that natural images simultaneously encompass rich low- and high-frequency components (Chan & Shen, 2005). Therefore, single natural image fitting has become an ideal test bed for INR models (Sitzmann et al., 2020; Saragadam et al., 2023; Xie et al., 2023; Shi et al., 2024a).
@@
In this experiment, we attempt to parameterize the function $\phi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3},\mathbf{x}\rightarrow\phi(\mathbf{x})$ that represents a discrete image. Following the previous work (Saragadam et al., 2023), we establish four-layer MLPs with $256$ hidden features. We conduct experiments on three MLPs architectures, i.e., MLPs with ReLU activation function (ReLU), ReLU with Positional encoding (PE) (Tancik et al., 2020) and MLPs with periodic activation function Sine (SIREN) (Sitzmann et al., 2020). These three models are classic baselines in the field of INRs and are widely used for comparison (Sitzmann et al., 2020; Saragadam et al., 2023; Shi et al., 2024a). We test on the first 8 images from the Kodak 24 dataset, each with a resolution of 768 × 512 pixels. For these images, ${\bm{K}}$ has approximately $10^{11}$ entries, rendering decomposition and multiplication infeasible during training. For our IGA method, we partition each image into non-overlapping $32\times 32$ patches as groups and points with the largest residuals are sampled to form the sample set ${\bm{X}}_{e}$. We construct the corresponding transformation matrix $\tilde{{\bm{S}}}_{e}$ with $end=20$ for SIREN and PE and $end=25$ for ReLU due to its severe spectral bias. Therefore, we generalize the inductive gradient adjustments of less than $0.1\%$ of total data points. For all architectures, we train baseline with vanilla gradients and with our inductive gradient adjustments (+IGA) by Adam optimizer. The learning rate schedule follows (Shi et al., 2024a), which maintains a fixed rate for the first $3000$ iterations and then reduces it by $0.1$ for another $7000$ iterations. We set initial learning rates as $35e-$ 3 for ReLU activation function and $31e-$ 3 for Sine activation function. For all baseline models, we set initial learning rates as $31e-$ 3 due to poor performances observed with $35e-$ 3. Current training adjustments methods, i.e., Fourier reparameterization (+FR) (Shi et al., 2024a) and batch normalization (+BN) (Cai et al., 2024) are also compared. Their hyperparameters follow publicly available codes, and we make every effort to achieve optimal performance. Full-batch training is adopted. In Table 1, we report three average metrics of different INRs over the first $8$ images of Kodak 24, where LPIPS values are measured by the “alex” from Zhang et al. (2018).
Figure 4: Visual examples of 2D color image approximation results by different training dynamics methods. Enlarged views of the regions labeled by red boxes are provided. The residuals of these regions in the Fourier domain are visualized through heatmaps in the top right corner. The increase in error corresponds to the transition of colors in the heatmaps from blue to red.
@@
As shown in Fig. 4, PE with our IGA method (PE+IGA) not only achieves the best improvements on PSNR values comparing to PE, PE+FR and PE+BN, but also facilitates a more precise representation of high-frequency details instead of being overly smooth like other methods. Concretely, as shown in the spectra located in the top right corner of Fig. 4, improvements of IGA are uniformly distributed across most frequency bands. In contrast, while BN and FR increase PSNR values, most improvements are near the origin, i.e., in the low-frequency range. This observation indicates that our IGA method allows for more balanced convergence rates of a wide spectral range with millions of data points. More visualization can be found in our Appendix D.
Table 1: Average metrics of 2D color image approximation results by different methods. The detailed settings can be found in Sec. 6.1. Per-image results are provided in our Appendix D.
Average ReLU PE SIREN
Metric Vanilla +FR +BN +IGA Vanilla +FR +BN +IGA Vanilla +FR +BN +IGA
PSNR $\uparrow$ 21.78 22.14 22.55 23.00 28.64 29.74 31.65 32.46 32.65 32.61 32.35 33.48
SSIM $\uparrow$ 0.483 0.492 0.505 0.513 0.783 0.817 0.870 0.882 0.898 0.899 0.894 0.912
LPIPS $\downarrow$ 0.630 0.631 0.595 0.555 0.222 0.187 0.114 0.090 0.081 0.081 0.102 0.067
@@
3D shape representation by Signed Distanced Functions (SDFs) has the advantage of modeling complex topologies. Therefore, representing SDFs has been widely adopted to test INRs models (Sitzmann et al., 2020; Saragadam et al., 2023; Cai et al., 2024; Shi et al., 2024a). In this section, we evaluate IGA on this task. Five 3D objects from the public dataset (Martel et al., 2021; Zhu et al., 2024) are utilized by previous settings of Saragadam et al. (2023); Shi et al. (2024a), which sample data points over a $512^{3}$ grid. We unfold these data points dimensionally and partition them into numerous groups, each containing $256$ points. For each iteration, we randomly draw $512$ groups and generalize the inductive gradient adjustments on these groups. The sampling strategy and MLP architecture are consistent with Experiment 6.1. We use Adam optimizer to minimize the $\ell_{2}$ loss between voxel values and INRs approximations. For fair comparison, the same training strategy is adopted for FR, BN and IGA. The detailed training strategy can be found in our Appendix E.
Table 2: Intersection over Union (IOU) of 3D shape representation by different methods. The detailed settings can be found in Sec. 6.2. The results for each scenario are provided in Appendix E.
Average ReLU PE SIREN
Metric Vanilla +FR +BN +IGA Vanilla +FR +BN +IGA Vanilla +FR +BN +IGA
IOU $\uparrow$ 9.647e-1 9.654e-1 9.542e-1 9.733e-1 9.942e-1 9.961e-2 9.938e-1 9.970e-1 9.889e-1 9.866e-1 9.825e-1 9.897e-1
Figure 5: Visual examples of 3D shape representation results by different training dynamics methods. Five images on the right correspond to the enlarged views of the red-boxed area of five models. More results can be found in our Appendix E.
In Table 2, we report the average intersection over union (IOU) metrics of five objects for reference. Under our training settings, baseline models optimized by vanilla gradient have converged to a favorable optimum, significantly outperforming prior works (Saragadam et al., 2023; Cai et al., 2024; Shi et al., 2024a). Nevertheless, our IGA method enables models to explore superior optima by purposefully improving the spectral bias, thereby leading to further improvements in representation accuracy such as more sharper edges in Fig. 5. More results can be found in our Appendix E.

### v2

@@ 4 Empirical Analysis on Simple Function Approximation
In this section, we aim to explore and provide empirical evidence for the theoretical results presented in Sec. 3.3. In experiment 1, we investigate the effects of ${\bm{K}}$-based and $\tilde{{\bm{K}}}$-based adjustments on spectral bias and observe that their differences diminish as the network width $m$ increases (Theorem 3.1). Training dynamics curves further suggest that increasing width $m$ reduces estimate error $\epsilon_{1}+\epsilon_{2}$ from inductive generalization (Theorem 3.2). In experiment 2, we validate $\tilde{{\bm{K}}}$-based adjustments and IGA on more general settings, demonstrating that IGA purposefully improves spectral bias, with its impacts controllable by managing the number of balanced eigenvalues in Eq. 5. This suggests the effectiveness of inductive generalization in estimating the training dynamics (Theorem 3.2) on general settings.
Figure 5: Visual examples of 3D shape representation results by different training dynamics methods. Five images on the right correspond to the enlarged views of the red-boxed area of ground truth and four models, respectively. More results can be found in our Appendix H.
Table 2: Intersection over Union (IOU) and Chamfer distance of 3D shape representation by different methods. The detailed settings can be found in Sec. 5.2. Per-scene results are provided in Table 12 and 13.
Average ReLU PE SIREN
Metric Vanilla +FR +BN +IGA Vanilla +FR +BN +IGA Vanilla +FR +IGA
IOU $\uparrow$ 9.647e-1 9.654e-1 9.662e-1 9.733e-1 9.942e-1 9.961e-2 9.946e-1 9.970e-1 9.889e-1 9.866e-1 9.897e-1
Chamfer Distance $\downarrow$ 5.936e-6 5.867e-6 5.610e-6 5.487e-6 5.123e-6 5.111e-6 5.116e-6 5.108e-6 5.688e-6 5.163e-6 5.157e-6
To better analyze, we follow Xu (2018); Shi et al. (2024a) and compute the relative error $\Delta_{k}$ between target signal $f$ and outputs $f_{\Theta}$ at frequency $k$ to show spectral bias:
$\Delta_{k}=|\mathcal{F}_{D}[f](k)-\mathcal{F}_{D}[f_{\Theta}](k)|/|\mathcal{F}_{D}[f](k)|,\vspace{-0.8em}$ (6)
@@
Experiment 1. In this experiment, as discussed above, we investigate the ${\bm{K}}$-based and $\tilde{{\bm{K}}}$-based adjustments. Therefore, the accurate computation of NTK matrix ${\bm{K}}$ is the key. Given that the accurate computation of ${\bm{K}}$ is non-trivial, we adopt a two-layer MLP with a fixed last layer as in Arora et al. (2019); Ronen et al. (2019), whose ${\bm{K}}_{ij}$ can be computed by the formula $\frac{1}{4\pi}(\mathbf{x}_{i}^{\top}\mathbf{x}_{j}+1)(\pi-\arccos(\mathbf{x}_{i}^{\top}\mathbf{x}_{j}))$. This architecture facilitates theoretical analysis but has limited representational capacity for complex functions. Considering these concerns, we constructed the following simple function $f(\cos(\theta),\sin(\theta)):\mathbb{S}^{1}\rightarrow\mathbb{R}^{1}$, widely used in prior works (Ronen et al., 2019; Geifman et al., 2023):
$\sin(0.4\pi\theta)+\sin(0.8\pi\theta)+\sin(1.6\pi\theta)+\sin(3.2\pi\theta).\vspace{-0.5em}$ (7)
@@
We set $N=1024$ for input samples via uniformly sampling $1024$ $\theta$ in $[0,2\pi]$. We vary the width of the MLP from $1024$ to $8192$. We adopt the Identity matrix ${\bm{I}}$ (that is, the vanilla gradient) and a series of ${\bm{S}},\tilde{{\bm{S}}},\tilde{{\bm{S}}}_{e}$ that $start=1$ and $end$ ranges from $10$ to $14$. Following our sampling strategy, we set $p=8$ and $n=128$, indicating that $\tilde{{\bm{S}}}_{e}$ is only $1/64$ the size of ${\bm{S}},\tilde{{\bm{S}}}$. All MLPs are trained separately with the same fixed learning rates by SGD for $20$K iterations. All models are repeated by 10 random seeds. We visualize the results of the baseline model and the adjustments with $end=14$ in Fig. 1. More results can be found in Appendix B.
@@
Analysis of Experiment 1. In Fig. 1, line plots show the convergence trends of MLPs by ${\bm{K}}$-based (${\bm{S}}$), $\tilde{{\bm{K}}}$-based ($\tilde{{\bm{S}}}$) and inductive gradient ($\tilde{{\bm{S}}}_{e}$) adjustments, compared to vanilla gradient (${\bm{I}}$). Heat maps with a consistent color scale visualize their spectral impacts, where darker colors indicate higher errors. Note that color bars of ${\bm{I}}$ darken rapidly from $0.4$Hz, highlighting severe spectral bias in MLPs trained with vanilla gradients. All the adjustments introduce lighter shades over 0.4Hz and 0.8Hz, effectively improving spectral bias, therefore resulting in faster MSE convergence. Please note that while adjustments by ${\bm{S}}$, $\tilde{{\bm{S}}}$ and $\tilde{{\bm{S}}}_{e}$ exhibit slight difference in MSE loss trends and impacts on spectrum in the case of width 1024, the differences are barely noticeable for width 4096. This result aligns with previous theoretical analysis that increasing width $m$ reduces differences between ${\bm{K}}$-based and $\tilde{{\bm{K}}}$-based adjustments (Theorem 3.1) and errors induced by inductive generalization (Theorem 3.2). Moreover, despite the presence of difference, $\tilde{{\bm{S}}}$ and $\tilde{{\bm{S}}}_{e}$ effectively improve spectral bias and the approximation accuracy, indicating that $\tilde{{\bm{K}}}$ successfully links spectral bias with training dynamics and our IGA method is effective.
@@
Experiment 2. In this experiment, we aim to validate $\tilde{{\bm{K}}}$-based adjustments and our IGA on more general settings. For IGA, we also aim to show that impacts on spectral bias can be tailored by managing the number of balanced eigenvalues under varying $p$. Therefore, we consider two practical INR models, i.e., MLPs with ReLU (ReLU) and MLPs with Sine (SIREN) (Sitzmann et al., 2020) like Shi et al. (2024a). To analyze tailored impacts under varying $p$, we use settings in Rahaman et al. (2019) and construct a 1D function $f:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$ with abundant spectrum as follows:
$\vspace{-0.5em}\sin(20\pi x)$ $+\sin(40\pi x)+\sin(60\pi x)+\sin(80\pi x)$
$+\sin(100\pi x)+\sin(120\pi x).\vspace{-0.5em}$ (8)
@@
SIREN is trained to regress the $f(x)$ with $2048$ values uniformly sampled in $[-1,1]$. For ReLU, we halve the frequency of each component due to its limited representation capacity (Yüce et al., 2022). For IGA, we set $p$ to $1$, $4$ and $8$ and vary $end$ from $1$ to $7$ to construct the corresponding $\tilde{{\bm{S}}}_{e}$. When $p=1$, IGA degenerates to $\tilde{{\bm{K}}}$-based adjustment. We adopt a four-hidden-layer, $256$-width architecture for ReLU and SIREN, optimized by Adam for $10$K iterations with a fixed learning rate. All models are repeated by $10$ random seeds. In Fig. 2, we adopt heatmaps of relative error $\Delta_{k}$ evolution to show impacts on spectral bias. In Fig. 3, we visualize $\Delta_{k}$ curves for some frequencies to precisely demonstrate impacts of IGA on convergence rates.
@@
Analysis of Experiment 2. In Fig. 2, for $\tilde{{\bm{K}}}$-based adjustment and IGA, as more balanced eigenvalues are introduced, the proportion of lighter regions in heatmaps of both ReLU and SIREN grows, while the relative error curve decreases more rapidly in Fig. 3. These observations clearly illustrate that impacts on spectral bias by $\tilde{{\bm{K}}}$-based adjustment and our IGA can be amplified by increasing the number of balanced eigenvalues on general settings. Further, as shown in Fig. 3, although the relative error curves of ReLU exhibit subtle differences as $p$ increases, our IGA effectively amplifies impacts on spectral bias with more balanced eigenvalues. These results demonstrate the consistency between our theoretical strategy and practical outcomes, supporting the extension of IGA to more general applications.
@@ 5 Experiment on Vision Applications
In this section, we apply our inductive gradient adjustment (IGA) method to various vision applications of INRs. The superior and consistent improvements demonstrate the broad applicability and superiority of IGA.
@@
Single natural image fitting has become an ideal test bed for INR models (Saragadam et al., 2023; Xie et al., 2023; Shi et al., 2024a), as natural images contain both low- and high-frequency components (Chan & Shen, 2005). In this experiment, we attempt to parameterize the function $\phi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3},\mathbf{x}\rightarrow\phi(\mathbf{x})$ that represents a discrete image. Following Saragadam et al. (2023), we establish four-layer MLPs with $256$ neurons per hidden layer. We test three MLPs architectures, i.e., MLPs with ReLU activations (ReLU), ReLU with Positional Encoding (PE) (Tancik et al., 2020) and MLPs with periodic activation Sine (SIREN) (Sitzmann et al., 2020). These three models are classic baselines in INRs and are widely used for comparison (Sitzmann et al., 2020; Saragadam et al., 2023; Shi et al., 2024a). Results on more activations can be found in Appendix E. We test on the first 8 images from the Kodak 24 (Franzen, 1999), each containing 768 × 512 pixels. For these images, ${\bm{K}}$ has approximately $10^{11}$ entries, rendering decomposition and multiplication infeasible during training. Following our sampling strategy, we group each image into non-overlapping $32\times 32$ patches ($p=1024$) and sample data with the largest residuals to form the ${\bm{X}}_{e}$. We construct $\tilde{{\bm{S}}}_{e}$ with $end=20$ for SIREN and PE and $end=25$ for ReLU due to its severe spectral bias. Therefore, we generalize the inductive gradient adjustments, derived from less than $0.1\%$ of the sampled data, to the entire population. For all architectures, we train baseline with vanilla gradients and with our IGA (+IGA) by Adam optimizer. The learning rate schedule follows Shi et al. (2024a). Detailed training settings can be found in Appendix G. Current training dynamics methods, i.e., Fourier reparameterization (+FR) (Shi et al., 2024a) and batch normalization (+BN) (Cai et al., 2024), are compared. Their hyperparameters follow publicly available codes, and we make every effort to achieve optimal performance. In Table 1, we report average values of four metrics of different INRs over the first $8$ images of Kodak 24, where LPIPS values are measured by the “alex” from Zhang et al. (2018).
@@
As shown in Table 1 and Fig. 4, PE with our IGA (PE+IGA) not only achieves the best improvements on average approximation accuracy: PSNR values compared to PE, PE+FR, and PE+BN, but also excels in frequency-aware metrics: SSIM, MS-SSIM, with a more precise representation of high-frequency details instead of being overly smooth like other methods. Concretely, as shown in the spectra located in the top right corner of Fig. 4, improvements of IGA are uniformly distributed across most frequency bands. In contrast, while BN and FR increase PSNR values, most improvements are near the origin, i.e., in the low-frequency range. Furthermore, as shown in Table 8-11, IGA achieves consistent improvements across almost all scenarios and metrics. While FR and BN achieve competitive results in some scenarios, their performance is generally less stable and occasionally leads to suboptimal or negative outcomes. These observations indicate that our IGA method allows for more balanced convergence rates of a wide spectral range with millions of data. More results are in Appendix G.
@@
3D shape representation by Signed Distanced Functions (SDFs) is widely used for testing INR models (Saragadam et al., 2023; Cai et al., 2024; Shi et al., 2024a), due to its ability to model complex topologies. In this section, we evaluate IGA on this task using five 3D objects from the public dataset (Martel et al., 2021; Zhu et al., 2024) and objects are sampled over a $512^{3}$ grid following Saragadam et al. (2023); Shi et al. (2024a). For IGA, following our sampling strategy for high-dimensional data, we set $p=256$ and randomly draw $512$ groups and generalize the inductive gradient adjustments on these groups in each iteration. The architectures of MLPs are consistent with Experiment 5.1. We use the Adam optimizer to minimize the $\ell_{2}$ loss between voxel values and INRs approximations. For a fair comparison, the same training strategy is adopted for FR, BN and IGA. Detailed settings can be found in our Appendix H.
In Table 2, we report the average intersection over union (IOU) and Chamfer Distance metrics of five objects for reference. Under our training settings, baseline models optimized by vanilla gradient have converged to a favorable optimum, significantly outperforming prior works (Saragadam et al., 2023; Cai et al., 2024; Shi et al., 2024a). Nevertheless, our IGA method enables models to explore superior optima by purposefully improving the spectral bias, thereby leading to improvements in representation accuracy such as sharper edges in Fig. 5. More results are in our Appendix H.